An experiment was performed to compare the fracture toughness of high-purity 18 Ni maraging weel with commercial purity of th question b
Answer #1

Ho 2 s не - > s

pop 1 pop 2
sample mean x = 65.40 59.90
std deviation σ= 1.400 1.000
sample size n= 30 39
std error σx1-x2=√(σ21/n122/n2)    = 0.302
test stat z =(x1-x2o)/σx1-x2     = 1.66
p value : = 0.0485 (from excel:1*normsdist(-1.66)

fail to reject Ho ; the data does not suggest........

b)

Hypothesized mean difference Δo       = 5
true mean difference Δ                                                          = 6
first sample standard deviation =σ1    =   1.400
second sample std deviation =σ2= 1.000
first sample size =n1 = 30.000
second sample size =n2 = 39.000
standard error=(√(σ12/n1+σ12/n1))= 0.3016
for 0.001 level and right tail critival Zα= 3.090
rejecf Ho if x>= Δo +Zα*σx or x>= 5.932
P(Type II error) =P(Xbar<5.932| Δ=6)=P(Z<(5.932-6)/0.302)=P(Z<-0.23)=0.4090